If the K, of a monoprotic weak acid is 1.3 x 10 , what is the pH of a 0.27 M solution of this acid? pH =
Answer #1

Consider a mono protic weak acid 3 which dissociates in solution as: - HB - tt + Br The dissociation constant ka of this acidDissociation constant is always expresse as a ratio of product of concentration of the products to the concentration of the r& [tb] = cox ; where c is the initial concentration of the acid. In this case, c = 0.27 So can be written as: - Ka= (x)(x) =2 Thurs : - k +.C7) ____ Given Ka= 1:30 100 and c= 0.27m. se equation @ Desemes- - _ 2 - - 1. 3.X10 + N (I.3x10-6) + ५(1.2X16Thus [++ ] = x= o.sax 10-3 M The formula to calculate pt of an acid is pt - log[it] So we get pH = -log(0.59 x103) $42-43.229